ganeshie8 (ganeshie8):

numbers of form \( 1111111\ldots \)

OpenStudy (anonymous):

rare primes party xD

OpenStudy (kainui):

There was one question a while back that was like prove 343434343... are composite, I'd really like to figure that one out still, I don't know, I don't want to hijack this we can do that one some other time though maybe?

ganeshie8 (ganeshie8):

working on this proof : a) say \(n\) = number of \(1\)'s in that number. next consider the smallest such number whose prime factor is \(p \) prove that \(p-1\) is divisible by \(n\) for all \(p \ge 5\) b) using part a, find the smallest such number which is divisible by \(13\)

ganeshie8 (ganeshie8):

i gave up on that 34343... problem kai, its very hard to prove that with the elementary number theory tricks it seems

ganeshie8 (ganeshie8):

but this problem is fairly simple.. i dont have a proof yet though

OpenStudy (kainui):

Ok wait I was assuming this was base 2, but this is really base 10 right @ganeshie8 ?

ganeshie8 (ganeshie8):

Ahh yes yes...

ganeshie8 (ganeshie8):

In short : we need to find the smallest integer divisible by \(13\) of form \(111\ldots\)

OpenStudy (kainui):

Example: \[\Large 111=10^0+10^1+10^2 \\ \Large 111... = \sum_{k=0}^{n-1}10^k\]

OpenStudy (kainui):

Not sure if that helps us, but maybe

ganeshie8 (ganeshie8):

i don't think we can get a better start than that xD \[\large 111\ldots (\text{n times}) = \dfrac{10^n-1}{10-1}\]

OpenStudy (anonymous):

xD will i was like n mod 9 is the same as keep summing digits

OpenStudy (jhannybean):

Always has something to do with mods, haha.

ganeshie8 (ganeshie8):

looks there was a mistake in problem statement \(p=5\) doesn't make sense, this only works for \(p \color{Red}{\gt} 5\)

OpenStudy (anonymous):

ok i dont really get what the questions is like what do u mean by next consider the smallest such number whose prime factor is p

OpenStudy (anonymous):

like 111.....1=p^k ?

ganeshie8 (ganeshie8):

if possible im lookong for a proof with out mods and other fancy stuff @Jhannybean but most likely i wont get it ;)

OpenStudy (anonymous):

yeah what ever :P

ganeshie8 (ganeshie8):

yeah ANYTHING is fine for a first proof :) il give an example @Marki

OpenStudy (kainui):

Yeah I'm lost with @Marki on this one, that didn't make sense to me either haha

OpenStudy (jhannybean):

@Michele_Laino any ideas?

ganeshie8 (ganeshie8):

consider \(n = 8\) case : \[\large 11111111 = 73\times 152207\] \(73-1\) is divisible by \(8\) as desired also \(11111111 \) is the smallest number thats divisible by \(73\)

ganeshie8 (ganeshie8):

want another example ?

OpenStudy (anonymous):

seems nice ! yeah continue

ganeshie8 (ganeshie8):

basically that theorem helps in finding the prime factors of numbers of form \(111\ldots \) easily

OpenStudy (anonymous):

yep yep, i wonder how did u get it :O

ganeshie8 (ganeshie8):

idk how to prove it other than the trivial observation that these numbers are of form \(\dfrac{10^n-1}{9}\)

ganeshie8 (ganeshie8):

http://mathworld.wolfram.com/Repunit.html

OpenStudy (kainui):

@Marki I did that same thing earlier and laughed at myself XD

OpenStudy (anonymous):

\(\large \dfrac{10^{n+1}-1}{9}=(10-1)(10^{n-1}+2.10^{n-2}+3.10^{n-3}+.....+n )+(n+1)\)

OpenStudy (anonymous):

i remember it from a question u solved before @Kainui lol was a good time :3

OpenStudy (kainui):

Oh woah! Interesting apparently "Every prime repunit is a circular prime."

OpenStudy (kainui):

Oh wait no nevermind that's obvious and stupid... X_X

OpenStudy (anonymous):

now lets try what possibles we have

OpenStudy (anonymous):

circle in mod 9 xD but ppl here gone crazy when i speak of it i'll try something different

ganeshie8 (ganeshie8):

circular primes look interesting xD @mathmath333 so the number with 1038 ones is divisible by 13, but that need not be the smallest number right ?

OpenStudy (mathmath333):

u mean the 1038 is not the smallest one ?

OpenStudy (anonymous):

so let n+1 be our prime then \((10-1)(10^{n-1}+2.10^{n-2}+3.10^{n-3}+.....+n ) =1 \mod n\) hmm idk if this is work

ganeshie8 (ganeshie8):

https://www.wolframalpha.com/input/?i=Table%5B%2810%5En-1%29%2F9+mod+13%2C+%7Bn%2C5%2C20%7D%5D it turns out that the smallest number divisible by 13 is \((10^6-1)/9\)

OpenStudy (mathmath333):

by the 13 is a prime so u can apply the fermat u think

ganeshie8 (ganeshie8):

Yes!

OpenStudy (mathmath333):

thats why 13-1 should be the smallest

ganeshie8 (ganeshie8):

\[ \Large 111... = \sum_{k=0}^{n-1}10^k \equiv \dfrac{10^n-1}{9} \pmod {p}\]

OpenStudy (anonymous):

nice

ganeshie8 (ganeshie8):

i don't get it

ganeshie8 (ganeshie8):

Oh wait, since \(p \nmid 9\), we can use fermat theorem directly on \(\large 10^n-1 \equiv 0 \pmod {p}\)

ganeshie8 (ganeshie8):

are you saying \(n \) must be multiple of \(p-1\) for this to hold ? thats a very interesting observation, let me think a bit more...

OpenStudy (kainui):

``` n=1, 1 = 1 n=2, 11 = 1011 n=3, 111 = 1101111 n=4, 1111 = 10001010111 n=5, 11111 = 10101101100111 n=6, 111111 = 11011001000000111 n=7, 1111111 = 100001111010001000111 n=8, 11111111 = 101010011000101011000111 n=9, 111111111 = 110100111110110101111000111 n=10, 1111111111 = 1000010001110100011010111000111 n=11, 11111111111 = 1010010110010001100001100111000111 n=12, 111111111111 = 1100111011110101111010000000111000111 n=13, 1111111111111 = 10000001010110011011000100001000111000111 n=14, 11111111111111 = 10100001101100000001110101001011000111000111 n=15, 111111111111111 = 11001010000111000010010010011101111000111000111 n=16, 1111111111111111 = 11111100101000110010110111000101010111000111000111 n=17, 11111111111111111 = 100111011110010111111100100110110101100111000111000111 n=18, 111111111111111111 = 110001010101111101111011110000100011000000111000111000111 ``` Some data to play with I just made a Java program if you want to mess with it: ``` public class PlayingAround { public static void main(String... args) { for (long i = 1; i < 19; i++) { System.out.println("n="+ i + ", " +repunit(i) + " = " + Long.toBinaryString(repunit(i)) ); } } public static long repunit(long n) { return (pow(10, n) - 1) / 9; } public static long pow(long a, long b) { long c = 1; for (long i = 0; i < b; i++) { c *= a; } return c; } } ```

ganeshie8 (ganeshie8):

Oh you're converting it to binary

OpenStudy (kainui):

Oh yeah I forgot to mention I was interested in that and base 5, thanks for reminding me lol.

OpenStudy (anonymous):

Hey you want to prove it or looking for some numbers?

ganeshie8 (ganeshie8):

its more or less proved by now @Catch.me :) let me put the proof in one reply quick

OpenStudy (kainui):

What's interesting me is the last digits converge to a repeating pattern ...000111000111 so I'm trying to see if there's some kind of pattern we can take from that to apply here.

ganeshie8 (ganeshie8):

By fermat's little thm we have (suggested by @mathmath333 earlier) \[10^{p-1}-1\equiv 0 \pmod {p}\] and we're searching for a prime \(p\) that satisfies \[10^{n}-1\equiv 0 \pmod {p}\] combining both we can conclude \(\large n | (p-1)\) \(\blacksquare \)

OpenStudy (mathmath333):

by fermat we can say it is true for table of 12 ... but why it true for table of 6 though

OpenStudy (mathmath333):

although table of 6 will cover the table of 12..

ganeshie8 (ganeshie8):

last digits for these reptunits in binary are \(\ldots 111000111000111\) when n = 2,19,23 we get prime repunits @Kainui could you generate a few more numbers xD

ganeshie8 (ganeshie8):

there cannot be any special pattern for primes just by changing the base.. but just wondering how the prime repunits look in binary overall

ganeshie8 (ganeshie8):

@mathmath333 once we find that \( n | (13-1)\), we keep testing all the factors of \(12\)

ganeshie8 (ganeshie8):

factors of 12 are {1,2,3,4,6,12} By testing them from least factor, we can conclude n=6 gives the smallest number divisible by 13

OpenStudy (mathmath333):

i was wondering it generates a new problem to prove \(\large\tt \begin{align} \color{black}{ \dfrac{10^{6n}-1}{9} (mod~13) \equiv 0 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

did i miss too much xD

ganeshie8 (ganeshie8):

are you suggesting every 6th repunit is a multiple of 13 ?

ganeshie8 (ganeshie8):

that looks pretty interesting xD

OpenStudy (mathmath333):

yes :)

OpenStudy (anonymous):

i was like amazed by this \(\Large 111... = \sum_{k=0}^{n-1}10^k \equiv \dfrac{10^n-1}{9} \pmod {p} \) now let n|p-1 then by fermat this holds \(\Large 111... = \sum_{k=0}^{n-1}10^k \equiv \dfrac{10^{p-1} -1}{9} \pmod {p} \)

ganeshie8 (ganeshie8):

yes i feel bit down for not thinking that lol, i had no clue about it till @mathmath333 suggested >.<

OpenStudy (anonymous):

the funny thing im trying to review primitive roots and it dint come into my mind :P

OpenStudy (mathmath333):

it was obvious by looking 13 which is prime

OpenStudy (anonymous):

xD im idiot lol

OpenStudy (anonymous):

ok so now what are u trying to find ?

OpenStudy (kainui):

I think I fell asleep at my keyboard, I'm out! Try wolfram alpha for more binaries, unfortunately the long data type breaks down after that point, not enough digits in memory. You need to use BigInteger or something I think to do that. Ok goodnight, have fun. XD

OpenStudy (anonymous):

why binary :O what have i missed :(

OpenStudy (kainui):

Maybe we should try working on project euler questions some time? This might interest you https://projecteuler.net/problem=35

ganeshie8 (ganeshie8):

looks we have proved more than what the problem is asking for : can we conclude if \(R_n\) is the least repunit number thats divisible by \(p \gt 5\), then every \(nk\)'th repunit is also divisible by \(p\) ? (\(k \in \mathbb{N}\))

OpenStudy (anonymous):

yep indeed!

ganeshie8 (ganeshie8):

evidence : https://www.wolframalpha.com/input/?i=Table%5B%2810%5E%286n%29-1%29%2F9+mod+13%2C+%7Bn%2C1%2C100%7D%5D im still bit unsure if it works in general though

ganeshie8 (ganeshie8):

@Kainui that question looks more like programming xD i remember working it with @dan815 couple of years ago

OpenStudy (anonymous):

ok since n|p-1 from fermat 10^p-1 =1 mod p so this hold (n is order of 10 from pattern ) 10^n=1 mod p \(\Large \dfrac{10^n-1}{9}=0 \pmod {p}\) thus \(\Large \dfrac{10^nk-1}{9}=0 \pmod {p}\)

OpenStudy (anonymous):

\(\Large \dfrac{10^{nk}-1}{9}=0 \pmod {p}\)*

ganeshie8 (ganeshie8):

how do you know \(n\) is the order of \(10\) for all primes ?

OpenStudy (anonymous):

from pattern we got( order only for p's that the pattern divide) \(\Large 111... = \sum_{k=0}^{n-1}10^k \equiv \dfrac{10^n-1}{9} \pmod {p}\) right ? gcd(9,p)=1 since p>5 thus \( 10^n-1 =0 \pmod {p}\) \( 10^n =1 \pmod {p}\)

OpenStudy (anonymous):

i might been late xD

ganeshie8 (ganeshie8):

yes \(10^n \equiv 1 \pmod p \) part is clear how do you know there is not true for any other integer less than \(n\) ?

ganeshie8 (ganeshie8):

thats the definition of order right ?

OpenStudy (anonymous):

any other integer less than n have the order let be k then k=vn done :P

OpenStudy (anonymous):

yeah but u got what im saying right ?

ganeshie8 (ganeshie8):

not yet, let me read again

ganeshie8 (ganeshie8):

order of \(a\) in \(\mod~n\) is the smallest positive integer \(k\) satisfying \(a^k \equiv 1\pmod n\)

OpenStudy (anonymous):

\( a^k\equiv a^v \equiv 1\pmod n \) and k<v then k|v

ganeshie8 (ganeshie8):

OMG! okay i got it xD

ganeshie8 (ganeshie8):

there cannot be any integers less than \(n\) satisfying the given congruence because \(R_n\) is the smallest repunit thats divisible by \(p\) that makes \(n\), the order of 10 in mod p pretty neat xD

OpenStudy (anonymous):

Very Interesting

OpenStudy (anonymous):

that gives me an idea for our main lovely problem

OpenStudy (anonymous):

oh ur enjoying it @marylou004